OpenStudy (anonymous):

simply the square root of 4x^2y^4

OpenStudy (anonymous):

2xy

OpenStudy (fibonaccichick666):

So, first can you simplify \(\sqrt(x^2)\)

OpenStudy (fibonaccichick666):

How about \(\sqrt(x^4)\)?

OpenStudy (anonymous):

Be careful \[ \sqrt { x^2} = |x| \]

OpenStudy (fibonaccichick666):

actually, it's + or -

OpenStudy (anonymous):

\[ \sqrt{(-5)^2}=5 =|-5| \]

OpenStudy (fibonaccichick666):

Absolute value is not required technically in front of every square root, there is a + or -

OpenStudy (anonymous):

No sir

OpenStudy (fibonaccichick666):

Yes, there is.

OpenStudy (fibonaccichick666):

Look at the quadratic formula for reference

OpenStudy (anonymous):

The square root is always positive or zero

OpenStudy (fibonaccichick666):

Noooooo, How about \(\sqrt(-4)\)? That is neiither

OpenStudy (fibonaccichick666):

A square root always has at least 2 possible answers provided the expression under the radical is not 0

OpenStudy (fibonaccichick666):

Although, I would prefer to ask these to the poser of the question and not argue semantics on a basic problem.

OpenStudy (anonymous):

I meant the function f(x) defined by \[ f(x) = \sqrt x \] Has only one value which is the positive one. Otherwise it will not be a function.

OpenStudy (anonymous):

Have Fun

OpenStudy (fibonaccichick666):

Yes, in terms of functions, I agree. Thank you. You as well.

OpenStudy (anonymous):

For me \[ \sqrt{ 4x^2y^4}= 2 |x| y^2 \]

OpenStudy (fibonaccichick666):

(function was not an initial condition here) Instead of the abs value I would put \[\left(\begin{matrix}+ \\ -\end{matrix}\right)2xy^2\] but like I said it's semantics in the long run.

OpenStudy (fibonaccichick666):

We are arguing over the same answer, just written differently. A different approach

OpenStudy (anonymous):

\[ \sqrt{ 4(-3)^2(2)^4}= 2 |-3| 2^2=2(3)(4)=24 \]

OpenStudy (fibonaccichick666):

In that case, I disagree with your result.

OpenStudy (anonymous):

The + is for x >0 The - for x <0 The |x| takes care of this distinction

OpenStudy (fibonaccichick666):

I always keep the distinction

OpenStudy (fibonaccichick666):

In this case I would have a -24. Due to the initial being a -3.

OpenStudy (fibonaccichick666):

But nonetheless, I think it is a different approach. I have never seen your approach, though, I do understand why one would take it.